# Privacy-Preserving Retrieval Augmented Generation with Differential Privacy

### v1

@@ 5 Related Work
[40] proposed an empirical privacy-preserving algorithm for RAG through the synthetic data generation, while our work studies privacy-preserving RAG in the framework differential privacy, which protects the privacy of each individual documents with the theoretical guarantee. Differential privacy has been studied in many other tasks in large language models too. Prompt tuning helps tailor the LLM to new tasks from a (private) test-domain dataset. [15] and [7] study the DP mechanism on two different prompt tuning frameworks: prompt optimization and offsite prompt tuning [34]. In-context learning adapts to different tasks by illustrating some examples in the context as the task description. DP in-context learning consider the situation when the examples are picked from any private set. [35] tackles this problem by generating synthetic examples with DP and [37] solves the DP test-query by generating the answers, both in a sample-and-aggregate fashion. The differentially private pretraining and finetuning of LLMs has been studied to address the privacy concern in the training data and memory is a large bottleneck when naively deploying DP-SGD [1]. [24] focuses on the pretraining stage which introduces ghost clipping to make DP-SGD more memory efficient. [38] explores finetuning in the parameter-efficient framework LoRA [17]. Notice that DP voting plays a crucial role in these sample-and-aggregate algorithms, including ours. A basic approach is to apply the Laplacian or Gaussian mechanism [12]. [27, 28] proposed a data-dependent privacy analysis, which can be tighter when the majority vote has a large margin over other options. We integrate the LimitedDomain mechanism for our algorithm, which addresses challenges when the voting domain is large [8]; the large vocabulary size in token voting is our main bottleneck.
@@
Our algorithms generate the answers token by token, where each token needs a query to the private dataset and consumes some privacy budget. In this paper, we set up the privacy parameters before the start of the algorithm and have a pre-set maximum number of token to generate. However, the number of tokens to generate is different per question and is unknown before the algorithm starts – it is possible that the number of generated tokens is much smaller than the pre-set number but we still need to pay the full pre-defined privacy cost. A line of work [32, 14, 22, 36] tries to measure the privacy budget in fully adaptive composition where the budget consuming can interact with the data. Especially, [36] gives an analysis for this fully adaptive setting which matches the tightness of advanced composition. The idea of fully adaptive composition sounds a fit to our problem, which allows us to “pay as we go”, rather than predefining the $\varepsilon_{\max}$ before the generation process. We found the analysis for fully adaptive setting is effective for large number of steps and small budget per step, while in our algorithm the number of generated tokens would not be very large and each token generation needs a relatively large budget to guarantee the utility. This mismatch makes us stick with the advanced composition.
@@ 6 Conclusion and Future Work
We introduce the first differentially private algorithms for RAG, enabling us to enhance LLMs by domain-specific but sensitive external corpus. With our novel combination of the DP voting algorithm and sparse vector technique along with the non-private LLM, we succeed in spending privacy budget only when the LLM needs sensitive information to generate a new token. Consequently, DPSparseVoteRAG successfully generates a sufficiently long and accurate response under a reasonable privacy budget. Our experimental results demonstrate that our algorithms perform better than the non-RAG baseline across different datasets and models, showing their effectiveness.
One of our future directions is to conduct more practical empirical evaluations. The Wikipedia dataset, which we use as the external data source, is typically included in the training data of recent LLMs. RAG is particularly effective when the external knowledge is truly sensitive and thus outside the LLM training data. It is essential to conduct evaluations that are as close to the real situation. Since our usage of the sparse vector technique is applicable to any DP token generation algorithm through voting, another future direction would be to examine how it improves DP token generation across different tasks, e.g., in-context learning and prompt tuning.
@@ Acknowledgments
Appendix A Additional Experimental Results
A.1 Main Results across Models and Datasets
In Figure 4, we present the average match accuracy of baseline algorithms and our algorithms for different total privacy guarantees ($\epsilon_{\mathrm{total}}$) with OPT (2.7B). We see the similar trend observed in Figure 3.
Figure 4: Average match accuracy comparison across algorithms on Trivia (left) and NQ (right) datasets with OPT (2.7B). We report the best results over hyperparameters for each $\epsilon_{\mathrm{total}}$.
For completeness, we further present the detailed results, in the same form of the one provided in Table 1, with OPT (1.3B and 2.7B), GPT2-XL and Pythia (1.4B) on Trivia and NQ datasets in Tables 28.
Table 2: Average match accuracy comparison of our algorithms on NQ dataset with OPT (1.3B) under varying values of total privacy budget $\epsilon_{\mathrm{total}}$ with different hyperparameters, the number of voters $m$ and $\epsilon_{\mathrm{token}}$. Bold font values represent the best performance of each algorithm under fixed $\epsilon_{\mathrm{total}}$.
@@ A.1 Main Results across Models and Datasets
$\epsilon_{\mathrm{token}}=1$ 0.20 0.21 0.23 0.27 0.32 0.32 0.27 0.32 0.32 0.27 0.32 0.32
$\epsilon_{\mathrm{token}}=2$ 0.03 0.03 0.03 0.26 0.23 0.23 0.36 0.35 0.36 0.36 0.35 0.36
$\epsilon_{\mathrm{token}}=5$ 0.00 0.00 0.00 0.03 0.03 0.03 0.24 0.22 0.22 0.33 0.31 0.30
@@
$\epsilon_{\mathrm{token}}=1$ 0.11 0.20 0.28 0.11 0.21 0.28 0.11 0.21 0.28 0.11 0.21 0.28
$\epsilon_{\mathrm{token}}=2$ 0.21 0.24 0.25 0.25 0.36 0.41 0.27 0.37 0.42 0.27 0.37 0.42
$\epsilon_{\mathrm{token}}=5$ 0.07 0.07 0.07 0.26 0.25 0.26 0.40 0.35 0.37 0.46 0.43 0.47
Table 3: Average match accuracy comparison of our algorithms on Trivia dataset with OPT (2.7B) under varying values of total privacy budget $\epsilon_{\mathrm{total}}$ with different hyperparameters, the number of voters $m$ and $\epsilon_{\mathrm{token}}$. Bold font values represent the best performance of each algorithm under fixed $\epsilon_{\mathrm{total}}$.
@@
$\epsilon_{\mathrm{token}}=1$ 0.34 0.38 0.38 0.44 0.48 0.51 0.44 0.48 0.51 0.44 0.48 0.51
$\epsilon_{\mathrm{token}}=2$ 0.03 0.03 0.03 0.39 0.40 0.40 0.50 0.52 0.54 0.50 0.52 0.54
$\epsilon_{\mathrm{token}}=5$ 0.00 0.00 0.00 0.03 0.03 0.03 0.33 0.30 0.28 0.48 0.48 0.48
@@
$\epsilon_{\mathrm{token}}=1$ 0.36 0.48 0.57 0.36 0.48 0.57 0.36 0.48 0.57 0.36 0.48 0.57
$\epsilon_{\mathrm{token}}=2$ 0.45 0.51 0.52 0.58 0.63 0.68 0.60 0.64 0.70 0.60 0.64 0.70
$\epsilon_{\mathrm{token}}=5$ 0.28 0.34 0.38 0.49 0.54 0.55 0.60 0.65 0.66 0.63 0.69 0.71
Table 4: Average match accuracy comparison of our algorithms on NQ dataset with OPT (2.7B) under varying values of total privacy budget $\epsilon_{\mathrm{total}}$ with different hyperparameters, the number of voters $m$ and $\epsilon_{\mathrm{token}}$. Bold font values represent the best performance of each algorithm under fixed $\epsilon_{\mathrm{total}}$.
@@
$\epsilon_{\mathrm{token}}=1$ 0.24 0.29 0.32 0.34 0.42 0.45 0.34 0.42 0.45 0.34 0.42 0.45
$\epsilon_{\mathrm{token}}=2$ 0.03 0.03 0.03 0.32 0.32 0.32 0.44 0.47 0.46 0.44 0.47 0.46
$\epsilon_{\mathrm{token}}=5$ 0.00 0.00 0.00 0.03 0.03 0.03 0.30 0.30 0.30 0.37 0.36 0.37
@@
$\epsilon_{\mathrm{token}}=1$ 0.15 0.18 0.24 0.15 0.19 0.25 0.15 0.19 0.25 0.15 0.19 0.25
$\epsilon_{\mathrm{token}}=2$ 0.25 0.29 0.33 0.36 0.46 0.52 0.37 0.48 0.54 0.37 0.48 0.54
$\epsilon_{\mathrm{token}}=5$ 0.11 0.09 0.10 0.33 0.35 0.34 0.51 0.52 0.50 0.53 0.56 0.57
Table 5: Average match accuracy comparison of our algorithms on Trivia dataset with GPT2-XL under varying values of total privacy budget $\epsilon_{\mathrm{total}}$ with different hyperparameters, the number of voters $m$ and $\epsilon_{\mathrm{token}}$. Bold font values represent the best performance of each algorithm under fixed $\epsilon_{\mathrm{total}}$.
@@
$\epsilon_{\mathrm{token}}=1$ 0.29 0.33 0.34 0.32 0.39 0.42 0.32 0.39 0.42 0.32 0.39 0.42
$\epsilon_{\mathrm{token}}=2$ 0.16 0.16 0.16 0.37 0.36 0.35 0.45 0.44 0.43 0.45 0.44 0.43
$\epsilon_{\mathrm{token}}=5$ 0.03 0.03 0.03 0.17 0.16 0.17 0.34 0.31 0.32 0.43 0.41 0.42
@@
$\epsilon_{\mathrm{token}}=1$ 0.32 0.35 0.37 0.32 0.35 0.37 0.32 0.35 0.37 0.32 0.35 0.37
$\epsilon_{\mathrm{token}}=2$ 0.37 0.34 0.42 0.42 0.44 0.54 0.43 0.46 0.56 0.43 0.46 0.56
$\epsilon_{\mathrm{token}}=5$ 0.24 0.25 0.24 0.41 0.41 0.42 0.54 0.51 0.52 0.57 0.56 0.58
Table 6: Average match accuracy comparison of our algorithms on NQ dataset with GPT2-XL under varying values of total privacy budget $\epsilon_{\mathrm{total}}$ with different hyperparameters, the number of voters $m$ and $\epsilon_{\mathrm{token}}$. Bold font values represent the best performance of each algorithm under fixed $\epsilon_{\mathrm{total}}$.
@@
$\epsilon_{\mathrm{token}}=1$ 0.20 0.30 0.30 0.23 0.33 0.34 0.23 0.33 0.34 0.23 0.33 0.34
$\epsilon_{\mathrm{token}}=2$ 0.14 0.14 0.15 0.33 0.33 0.32 0.37 0.37 0.36 0.37 0.37 0.36
$\epsilon_{\mathrm{token}}=5$ 0.04 0.04 0.05 0.13 0.14 0.16 0.29 0.30 0.31 0.35 0.37 0.36
@@
$\epsilon_{\mathrm{token}}=1$ 0.09 0.19 0.19 0.09 0.19 0.20 0.09 0.19 0.20 0.09 0.19 0.20
$\epsilon_{\mathrm{token}}=2$ 0.24 0.26 0.32 0.27 0.34 0.36 0.28 0.36 0.38 0.28 0.36 0.38
$\epsilon_{\mathrm{token}}=5$ 0.11 0.12 0.09 0.32 0.31 0.32 0.39 0.41 0.39 0.42 0.43 0.42
Table 7: Average match accuracy comparison of our algorithms on Trivia dataset with Pythia (1.4B) under varying values of total privacy budget $\epsilon_{\mathrm{total}}$ with different hyperparameters, the number of voters $m$ and $\epsilon_{\mathrm{token}}$. Bold font values represent the best performance of each algorithm under fixed $\epsilon_{\mathrm{total}}$.
@@
$\epsilon_{\mathrm{token}}=1$ 0.23 0.26 0.27 0.34 0.39 0.41 0.34 0.39 0.41 0.34 0.39 0.41
$\epsilon_{\mathrm{token}}=2$ 0.14 0.14 0.14 0.29 0.30 0.30 0.43 0.46 0.44 0.43 0.46 0.44
$\epsilon_{\mathrm{token}}=5$ 0.03 0.03 0.03 0.14 0.14 0.14 0.27 0.27 0.27 0.38 0.39 0.38
@@
$\epsilon_{\mathrm{token}}=1$ 0.26 0.31 0.39 0.26 0.31 0.39 0.26 0.31 0.39 0.26 0.31 0.39
$\epsilon_{\mathrm{token}}=2$ 0.38 0.37 0.40 0.45 0.47 0.48 0.46 0.52 0.53 0.46 0.52 0.53
$\epsilon_{\mathrm{token}}=5$ 0.22 0.23 0.24 0.40 0.39 0.38 0.52 0.52 0.51 0.56 0.57 0.56
Table 8: Average match accuracy comparison of our algorithms on NQ dataset with Pythia (1.4B) under varying values of total privacy budget $\epsilon_{\mathrm{total}}$ with different hyperparameters, the number of voters $m$ and $\epsilon_{\mathrm{token}}$. Bold font values represent the best performance of each algorithm under fixed $\epsilon_{\mathrm{total}}$.
@@
$\epsilon_{\mathrm{token}}=1$ 0.16 0.19 0.19 0.20 0.24 0.25 0.20 0.24 0.25 0.20 0.24 0.25
$\epsilon_{\mathrm{token}}=2$ 0.11 0.10 0.10 0.24 0.24 0.25 0.30 0.30 0.30 0.30 0.30 0.30
$\epsilon_{\mathrm{token}}=5$ 0.02 0.02 0.02 0.11 0.11 0.11 0.22 0.22 0.22 0.30 0.31 0.31

### v2

@@ 5 Related Work
[41] proposed an empirical privacy-preserving algorithm for RAG through the synthetic data generation, while our work studies privacy-preserving RAG in the framework differential privacy, which protects the privacy of each individual document with the theoretical guarantee. Differential privacy has been studied in many other tasks in large language models too. Prompt tuning helps tailor the LLM to new tasks from a (private) test-domain dataset. [16] and [7] study the DP mechanism on two different prompt tuning frameworks: prompt optimization and offsite prompt tuning [35]. In-context learning adapts to different tasks by illustrating some examples in the context as the task description. DP in-context learning considers the situation when the examples are picked from any private set. [36] tackles this problem by generating synthetic examples with DP and [38] solves the DP test query by generating the answers, both in a sample-and-aggregate fashion. The differentially private pretraining and finetuning of LLMs has been studied to address the privacy concern in the training data and memory is a large bottleneck when naively deploying DP-SGD [1]. [25] focuses on the pretraining stage which introduces ghost clipping to make DP-SGD more memory efficient. [39] explores finetuning in the parameter-efficient framework LoRA [18]. Notice that DP voting plays a crucial role in these sample-and-aggregate algorithms, including ours. A basic approach is to apply the Laplacian or Gaussian mechanism [13]. [28, 29] proposed a data-dependent privacy analysis, which can be tighter when the majority vote has a large margin over other options. We integrate the LimitedDomain mechanism for our algorithm, which addresses challenges when the voting domain is large [9]; the large vocabulary size in token voting is our main bottleneck.
@@
Our algorithms generate the answers token by token, where each token needs a query to the private dataset and consumes some privacy budget. In this paper, we set up the privacy parameters before the start of the algorithm and have a pre-set maximum number of tokens to generate. However, the number of tokens to generate is different per question and is unknown before the algorithm starts – it is possible that the number of generated tokens is much smaller than the pre-set number but we still need to pay the full pre-defined privacy cost. A line of work [33, 15, 23, 37] tries to measure the privacy budget in fully adaptive composition where the budget consuming can interact with the data. Especially, [37] gives an analysis for this fully adaptive setting which matches the tightness of advanced composition. The idea of fully adaptive composition sounds a fit to our problem, which allows us to “pay as we go”, rather than predefining the $\epsilon_{\mathrm{total}}$ before the generation process. We found the analysis for fully adaptive setting is effective for large number of steps and small budget per step, while in our algorithm the number of generated tokens would not be very large and each token generation needs a relatively large budget to guarantee the utility. This mismatch makes us stick with the advanced composition.
@@ 6 Conclusion and Future Work
We introduce the first differentially private algorithms for RAG, enabling us to enhance LLMs by domain-specific but sensitive external corpus. With our novel combination of the DP voting algorithm and sparse vector technique along with the non-private LLM, we succeed in spending privacy budget only when the LLM needs sensitive information to generate a new token. Consequently, DPSparseVoteRAG generates a sufficiently long and accurate response under a reasonable privacy budget. Our experiments demonstrate that our algorithms outperform the non-RAG baseline across different datasets and models, showing their effectiveness.
One of our future directions is to conduct more practical empirical evaluations. The Wikipedia dataset, which we use as the external data source, is typically included in the training data of recent LLMs. RAG is particularly effective when the external knowledge is truly sensitive and thus outside the LLM training data. It is essential for us to conduct evaluations that are as close to the real situation as possible and see how effective our algorithms are over non-RAG LLMs. Since our usage of the sparse vector technique is applicable to any DP token generation algorithm through voting, another future direction would be to examine how it improves DP token generation across different tasks, e.g., in-context learning and prompt tuning.
@@ Acknowledgments
Appendix A Proof of Theorem 1
Proof.
Let’s first consider steps 3 and 4 in Algorithm 1. Suppose $L_{x}$ is the list of documents in $D_{x}$ ranked by the relevance. One way to uniformly randomly split $D_{x}$ into $m$ disjoint subsets $D_{x}^{1},\cdots,D_{x}^{m}$ is that: given a ranked list of documents $L_{x}=(d_{1},\cdots,d_{mk})$, we randomly permute this list by $\pi$ to $L_{x}^{\pi}=(d_{\pi(1)},\cdots,d_{\pi(mk)})$ and let $D_{x}^{i}:=\{d_{\pi((i-1)k+1)},\cdots,d_{\pi(ik)}\}$. The process from $D_{x}$ to $L_{x}$ is deterministic, and the remaining of the algorithm is independent of $D$ given $L_{x}^{\pi}$. Therefore, we can equivalently denote the outcome of Algorithm 1 as $\mathcal{A}(L_{x}^{\pi})$.
For any two neighboring datasets $D$ and $D^{\prime}$, the retrieved datasets are $D_{x}=R(x,D;mk)$ and $D_{x}^{\prime}=R(x,D^{\prime};mk)$ and we denote $L_{x}=(d_{1},\cdots,d_{mk})$ and $L_{x}^{\prime}=(d_{1}^{\prime},\cdots,d_{mk}^{\prime})$. We only need to show for any set of outcomes $S$, $\Pr_{A,\pi}[A(L_{x}^{\pi})\in S]\leq\exp(\epsilon)\Pr_{A,\pi}[A((L_{x}^{\prime})^{\pi})\in S]+\delta.$ First of all, $D_{x}$ and $D_{x}^{\prime}$ have at most one different document (without considering the order). Therefore we can define another list of documents $L^{\prime\prime}_{x}=(d_{1}^{\prime\prime},\cdots,d_{mk}^{\prime\prime})$, such that $L^{\prime\prime}_{x}$ is some ranking of $D_{x}\textquoteright$ ’ and it differs at most one position from $L$. Notice that $L^{\prime}_{x}$ and $L^{\prime\prime}_{x}$ are only different at orders. Therefore $(L_{x}^{\prime})^{\pi}$ and $(L_{x}^{\prime\prime})^{\pi}$ have same distributions and as a consequence $A((L_{x}^{\prime})^{\pi})$ and $A((L_{x}^{\prime\prime})^{\pi})$ have same distributions.
Thus, the remaining is to prove for any set of outcomes $S$, $\Pr_{A,\pi}[A(L_{x}^{\pi})\in S]\leq\exp(\epsilon)\Pr_{A,\pi}[A((L_{x}^{\prime\prime})^{\pi})\in S]+\delta.$ We can actually prove a stronger conclusion $\Pr_{A}[A(L_{x}^{\pi})\in S]\leq\exp(\epsilon)\Pr_{A}[A((L_{x}^{\prime\prime})^{\pi})\in S]+\delta.$ It is because $L_{x}^{\pi}$ and $(L_{x}^{\prime\prime})^{\pi}$ differ at most one position and therefore at most one subset in step 4 is different given $L_{x}^{\pi}$ or $(L_{x}^{\prime\prime})^{\pi}$. This means that the histogram in step 10 differs at most one token. The guarantees of LimitedDomain and the composition theorem of DP together imply $\Pr_{A}[A(L_{x}^{\pi})\in S]\leq\exp(\epsilon)\Pr_{A}[A((L_{x}^{\prime\prime})^{\pi})\in S]+\delta.$
Appendix B Additional Experimental Results
B.1 Main Results across Models and Datasets
Figure 4: Average match accuracy comparison across algorithms on Trivia (left) and NQ (right) datasets with OPT (2.7B). The reported results are the means and standard deviations of average match accuracy over three runs. We report the best results over hyperparameters for each $\epsilon_{\mathrm{total}}$.
Figure 5: Average match accuracy comparison across algorithms on Trivia (left) and NQ (right) datasets with Llama 3.2 (1B). The reported results are the means and standard deviations of average match accuracy over three runs. We report the best results over hyperparameters for each $\epsilon_{\mathrm{total}}$.
Figure 6: Average match accuracy comparison across algorithms on Trivia (left) and NQ (right) datasets with GPT2-XL. The reported results are the means and standard deviations of average match accuracy over three runs. We report the best results over hyperparameters for each $\epsilon_{\mathrm{total}}$.
In Figures 46, we present the average match accuracy of baseline algorithms and our algorithms for different total privacy guarantees ($\epsilon_{\mathrm{total}}$) with OPT (2.7B), Llama 3.2 (1B), and GPT2-XL. We see the similar trend observed in Figure 3.
For completeness, we further present the detailed results, in the same form of the one provided in Table 1, with OPT (1.3B and 2.7B), Pythia (1.4B), Llama 3.1 (8B), Llama 3.2 (1B), and GPT2-XL on Trivia and NQ datasets in Tables 212.
Table 2: Average match accuracy comparison of our algorithms on NQ dataset with OPT (1.3B) under varying values of total privacy budget $\epsilon_{\mathrm{total}}$ with different hyperparameters, the number of voters $m$ and $\epsilon_{\mathrm{token}}$. Bold font values represent the best performance of each algorithm under fixed $\epsilon_{\mathrm{total}}$. We report the means of average match accuracy over three runs.
@@ B.1 Main Results across Models and Datasets
$\epsilon_{\mathrm{token}}=1$ $0.17$ $\bm{0.23}$ $\bm{0.23}$ $0.28$ $0.35$ $\bm{0.37}$ $0.28$ $0.35$ $0.37$ $0.28$ $0.35$ $0.37$
$\epsilon_{\mathrm{token}}=2$ $0.12$ $0.14$ $0.13$ $0.24$ $0.25$ $0.26$ $0.38$ $\bm{0.42}$ $\bm{0.42}$ $0.38$ $\bm{0.42}$ $\bm{0.42}$
$\epsilon_{\mathrm{token}}=5$ $0.04$ $0.06$ $0.05$ $0.12$ $0.14$ $0.13$ $0.22$ $0.22$ $0.21$ $0.36$ $0.36$ $0.35$
@@
$\epsilon_{\mathrm{token}}=1$ $0.14$ $0.18$ $0.28$ $0.14$ $0.18$ $0.29$ $0.14$ $0.18$ $0.29$ $0.14$ $0.18$ $0.29$
$\epsilon_{\mathrm{token}}=2$ $0.23$ $0.25$ $\bm{0.29}$ $0.37$ $0.41$ $\bm{0.43}$ $0.39$ $0.44$ $\bm{0.47}$ $0.39$ $0.44$ $0.47$
$\epsilon_{\mathrm{token}}=5$ $0.11$ $0.14$ $0.13$ $0.28$ $0.31$ $0.29$ $0.43$ $0.43$ $0.42$ $\bm{0.53}$ $\bm{0.53}$ $\bm{0.53}$
Table 3: Average match accuracy comparison of our algorithms on Trivia dataset with OPT (2.7B) under varying values of total privacy budget $\epsilon_{\mathrm{total}}$ with different hyperparameters, the number of voters $m$ and $\epsilon_{\mathrm{token}}$. Bold font values represent the best performance of each algorithm under fixed $\epsilon_{\mathrm{total}}$. We report the means of average match accuracy over three runs.
@@
$\epsilon_{\mathrm{token}}=1$ $0.34$ $\bm{0.40}$ $0.39$ $0.48$ $0.58$ $\bm{0.59}$ $0.48$ $0.58$ $0.59$ $0.48$ $0.58$ $0.59$
$\epsilon_{\mathrm{token}}=2$ $0.24$ $0.23$ $0.23$ $0.41$ $0.41$ $0.40$ $0.59$ $\bm{0.61}$ $0.60$ $0.59$ $\bm{0.61}$ $0.60$
$\epsilon_{\mathrm{token}}=5$ $0.09$ $0.10$ $0.10$ $0.23$ $0.24$ $0.23$ $0.38$ $0.39$ $0.38$ $0.49$ $0.51$ $0.51$
@@
$\epsilon_{\mathrm{token}}=1$ $0.31$ $0.48$ $\bm{0.53}$ $0.31$ $0.49$ $0.53$ $0.31$ $0.49$ $0.53$ $0.31$ $0.49$ $0.53$
$\epsilon_{\mathrm{token}}=2$ $0.50$ $0.52$ $\bm{0.53}$ $0.57$ $0.66$ $\bm{0.68}$ $0.58$ $0.69$ $\bm{0.70}$ $0.58$ $0.69$ $0.70$
$\epsilon_{\mathrm{token}}=5$ $0.35$ $0.38$ $0.37$ $0.54$ $0.54$ $0.54$ $0.66$ $0.66$ $0.66$ $\bm{0.73}$ $0.71$ $0.70$
Table 4: Average match accuracy comparison of our algorithms on NQ dataset with OPT (2.7B) under varying values of total privacy budget $\epsilon_{\mathrm{total}}$ with different hyperparameters, the number of voters $m$ and $\epsilon_{\mathrm{token}}$. Bold font values represent the best performance of each algorithm under fixed $\epsilon_{\mathrm{total}}$. We report the means of average match accuracy over three runs.
@@
$\epsilon_{\mathrm{token}}=1$ $0.26$ $0.30$ $\bm{0.31}$ $0.39$ $0.48$ $\bm{0.49}$ $0.39$ $0.48$ $0.49$ $0.39$ $0.48$ $0.49$
$\epsilon_{\mathrm{token}}=2$ $0.20$ $0.18$ $0.17$ $0.36$ $0.34$ $0.33$ $\bm{0.54}$ $0.51$ $0.51$ $\bm{0.54}$ $0.51$ $0.51$
$\epsilon_{\mathrm{token}}=5$ $0.06$ $0.05$ $0.05$ $0.20$ $0.17$ $0.17$ $0.34$ $0.31$ $0.30$ $0.46$ $0.43$ $0.42$
@@
$\epsilon_{\mathrm{token}}=1$ $0.16$ $0.24$ $0.32$ $0.16$ $0.24$ $0.32$ $0.16$ $0.24$ $0.32$ $0.16$ $0.24$ $0.32$
$\epsilon_{\mathrm{token}}=2$ $0.31$ $0.32$ $\bm{0.34}$ $0.46$ $0.52$ $\bm{0.54}$ $0.47$ $0.53$ $\bm{0.57}$ $0.47$ $0.53$ $0.57$
$\epsilon_{\mathrm{token}}=5$ $0.12$ $0.12$ $0.11$ $0.38$ $0.35$ $0.35$ $\bm{0.57}$ $0.53$ $0.53$ $\bm{0.61}$ $0.58$ $0.60$
Table 5: Average match accuracy comparison of our algorithms on Trivia dataset with Pythia (1.4B) under varying values of total privacy budget $\epsilon_{\mathrm{total}}$ with different hyperparameters, the number of voters $m$ and $\epsilon_{\mathrm{token}}$. Bold font values represent the best performance of each algorithm under fixed $\epsilon_{\mathrm{total}}$. We report the means of average match accuracy over three runs.
@@
$\epsilon_{\mathrm{token}}=1$ $0.24$ $0.28$ $\bm{0.29}$ $0.34$ $0.41$ $\bm{0.42}$ $0.34$ $0.41$ $0.42$ $0.34$ $0.41$ $0.42$
$\epsilon_{\mathrm{token}}=2$ $0.18$ $0.18$ $0.19$ $0.30$ $0.31$ $0.32$ $0.43$ $\bm{0.45}$ $0.44$ $0.43$ $\bm{0.45}$ $0.44$
$\epsilon_{\mathrm{token}}=5$ $0.06$ $0.05$ $0.07$ $0.18$ $0.18$ $0.19$ $0.29$ $0.30$ $0.30$ $0.38$ $0.39$ $0.37$
@@
$\epsilon_{\mathrm{token}}=1$ $0.25$ $0.32$ $0.38$ $0.25$ $0.32$ $0.39$ $0.25$ $0.32$ $0.39$ $0.25$ $0.32$ $0.39$
$\epsilon_{\mathrm{token}}=2$ $0.37$ $\bm{0.40}$ $\bm{0.40}$ $0.42$ $0.47$ $\bm{0.48}$ $0.43$ $0.53$ $\bm{0.54}$ $0.43$ $0.53$ $0.54$
$\epsilon_{\mathrm{token}}=5$ $0.25$ $0.25$ $0.26$ $0.43$ $0.44$ $0.41$ $0.51$ $0.53$ $0.51$ $\bm{0.57}$ $\bm{0.57}$ $\bm{0.57}$
Table 6: Average match accuracy comparison of our algorithms on NQ dataset with Pythia (1.4B) under varying values of total privacy budget $\epsilon_{\mathrm{total}}$ with different hyperparameters, the number of voters $m$ and $\epsilon_{\mathrm{token}}$. Bold font values represent the best performance of each algorithm under fixed $\epsilon_{\mathrm{total}}$. We report the means of average match accuracy over three runs.
@@
$\epsilon_{\mathrm{token}}=1$ $0.16$ $\bm{0.23}$ $\bm{0.23}$ $0.21$ $0.30$ $\bm{0.32}$ $0.21$ $0.30$ $0.32$ $0.21$ $0.30$ $0.32$
$\epsilon_{\mathrm{token}}=2$ $0.09$ $0.08$ $0.09$ $0.23$ $0.26$ $0.26$ $0.32$ $\bm{0.35}$ $\bm{0.35}$ $0.32$ $\bm{0.35}$ $\bm{0.35}$
$\epsilon_{\mathrm{token}}=5$ $0.05$ $0.04$ $0.04$ $0.10$ $0.09$ $0.09$ $0.23$ $0.24$ $0.23$ $0.31$ $0.32$ $0.30$
@@
$\epsilon_{\mathrm{token}}=1$ $0.09$ $0.17$ $0.24$ $0.09$ $0.17$ $0.24$ $0.09$ $0.17$ $0.24$ $0.09$ $0.17$ $0.24$
$\epsilon_{\mathrm{token}}=2$ $0.20$ $0.25$ $\bm{0.26}$ $0.25$ $0.36$ $\bm{0.41}$ $0.26$ $0.37$ $\bm{0.43}$ $0.26$ $0.37$ $0.43$
$\epsilon_{\mathrm{token}}=5$ $0.14$ $0.14$ $0.15$ $0.27$ $0.28$ $0.27$ $0.39$ $0.41$ $\bm{0.43}$ $0.46$ $\bm{0.51}$ $0.49$
Table 7: Average match accuracy comparison of our algorithms on Trivia dataset with Llama 3.1 (8B) under varying values of total privacy budget $\epsilon_{\mathrm{total}}$ with different hyperparameters, the number of voters $m$ and $\epsilon_{\mathrm{token}}$. Bold font values represent the best performance of each algorithm under fixed $\epsilon_{\mathrm{total}}$. We report the means of average match accuracy over three runs.
@@
$\epsilon_{\mathrm{token}}=1$ $0.75$ $0.77$ $\bm{0.79}$ $0.81$ $0.84$ $\bm{0.85}$ $0.81$ $0.84$ $\bm{0.85}$ $0.81$ $0.84$ $\bm{0.85}$
$\epsilon_{\mathrm{token}}=2$ $0.51$ $0.51$ $0.51$ $0.78$ $0.78$ $0.78$ $\bm{0.85}$ $\bm{0.85}$ $\bm{0.85}$ $\bm{0.85}$ $\bm{0.85}$ $\bm{0.85}$
$\epsilon_{\mathrm{token}}=5$ $0.20$ $0.21$ $0.21$ $0.51$ $0.51$ $0.51$ $0.73$ $0.74$ $0.74$ $0.83$ $0.82$ $0.82$
@@
$\epsilon_{\mathrm{token}}=1$ $0.72$ $0.83$ $0.88$ $0.72$ $0.83$ $0.88$ $0.72$ $0.83$ $0.88$ $0.72$ $0.83$ $0.88$
$\epsilon_{\mathrm{token}}=2$ $0.89$ $\bm{0.94}$ $\bm{0.94}$ $0.93$ $\bm{0.96}$ $\bm{0.96}$ $0.93$ $0.96$ $0.96$ $0.93$ $0.96$ $0.96$
$\epsilon_{\mathrm{token}}=5$ $0.76$ $0.78$ $0.76$ $0.93$ $0.93$ $0.93$ $0.95$ $\bm{0.97}$ $\bm{0.97}$ $0.95$ $\bm{0.97}$ $\bm{0.97}$
Table 8: Average match accuracy comparison of our algorithms on NQ dataset with Llama 3.1 (8B) under varying values of total privacy budget $\epsilon_{\mathrm{total}}$ with different hyperparameters, the number of voters $m$ and $\epsilon_{\mathrm{token}}$. Bold font values represent the best performance of each algorithm under fixed $\epsilon_{\mathrm{total}}$. We report the means of average match accuracy over three runs.
@@
$\epsilon_{\mathrm{token}}=1$ $0.49$ $\bm{0.54}$ $\bm{0.54}$ $0.57$ $0.63$ $\bm{0.64}$ $0.57$ $0.63$ $0.64$ $0.57$ $0.63$ $0.64$
$\epsilon_{\mathrm{token}}=2$ $0.30$ $0.29$ $0.30$ $0.55$ $0.56$ $0.55$ $0.64$ $\bm{0.66}$ $0.65$ $0.64$ $\bm{0.66}$ $0.65$
$\epsilon_{\mathrm{token}}=5$ $0.11$ $0.11$ $0.11$ $0.29$ $0.29$ $0.30$ $0.52$ $0.52$ $0.51$ $0.62$ $0.63$ $0.62$